- surrow's home page
- Posts
- 2024
- May (1)
- 2018
- 2017
- 2016
- 2015
- 2014
- October (3)
- September (1)
- August (1)
- July (1)
- June (3)
- May (4)
- April (6)
- March (4)
- February (5)
- January (2)
- 2013
- December (2)
- November (1)
- October (1)
- September (4)
- August (6)
- July (11)
- June (8)
- May (5)
- April (5)
- March (11)
- February (4)
- 2012
- December (3)
- November (3)
- October (3)
- September (3)
- August (3)
- July (10)
- June (9)
- May (5)
- April (4)
- March (4)
- February (7)
- January (4)
- 2011
- 2010
- December (2)
- November (4)
- August (2)
- July (3)
- June (6)
- May (4)
- April (4)
- March (8)
- February (6)
- January (4)
- 2009
- My blog
- Post new blog entry
- All blogs
FGT Commissioning Plan - Run 13
Commissioning Plan - FGT Run 13


(1) Documents:
(2) Setup:
Manpower at BNL: Xuan, Bernd, Gerrit, Akio, Ramiro, Maxence
Manpower outside BNL: Gerard, Anselm, Renee
Date: Monday, December 17, 2012 - Friday, February 15, 2013
Tasks:
(A) HV: Bernd, Xuan and Gerard
- Tests on HV modules with one broken channel: Friday, December 21, 2012
- Fix one broken HV channel on two modules (incl. spare module): Ship both modules during the week of Monday, December 24, 2012
- Reinstall and test repaired HV module, roughly during the week of Monday, January 21, 2013 (~ 4 weeks after shipment)
- Test spare HV module, roughly during the week of Monday, January 21, 2013 (~ 4 weeks after shipment)
- Modify labeling of slow control HV GUI by January 14, 2013
- Modify ramp profile and test on subset for HV modules 0/1 by January 21, 2013
- Full HV test after re-installation during the week of Monday, January 21, 2013
- Perform test ramp to nominal operating voltage of 3400V two days after switch to ArCO2, i.e. Wednesday, February 13, 2013
(B) DAQ: Gerard, Bernd and Xuan
* Complete DAQ system and verify proper working condition: Week of Monday, February 05, 2013
- Reinstall modules
- Pedestal runs
- Production of pedestal / status tables
- Pedestal runs with HV on / off
- APV parameter studies on pedestal / pedestal adjustment
* Deadline for providing complete ARC (I) (2) / ARM (12) Run 13 configuration: Thursday, January 31, 2013

* Deadline for providing spare ARC (I) (1) / ARM (2) FGT modules: Thursday, February 28, 2013
- Maintain GMT functionality
- Provide one ARC (I) module and two ARM modules by IST / IUCF groups to FGT group

(C) Gas: Gerrit, Bob
- Switch from N2 to ArCO2 (90/10) on Monday, February 11, 2013 provided that current schedule holds, i.e. roughly on week before overnight collisions.
(D) Documentation: Xuan, Bernd
- Update FGT shift-crew document by January 14, 2013
- Update gas system document by Don by January 14, 2013
- Update FGT Drupal page layout / documentation by January 14, 2013
- Update mapping documentation by January 14, 2013
(3) Commissioning:
Manpower at BNL: Xuan, Bernd, Akio, Ramiro, Maxence, Anselm, Devika
Manpower outside BNL: Gerard, Anselm, Renee
Date: February 15, 2013 - March 01, 2013
Fixed parameters / settings: ArCO2 (90/10) and nominal HV of 3400V / 7 time bins
(A) Overnight collision tests: February 15, 2013 - February 22, 2013
a) MinBias: 1 week
- Test DAQ system: Pedestal run
- Ramp HV under stable beam conditions: 3400V at 90/10 (C10)
- Timing scan of Run 12 latency / Establish collision signal
- Other APV parameter scans
- Re-adjustment of pedestal if necessary
- HV scan around 3400V
- Verify Quadrant mapping
- Verify APV mapping
- Verify Strip mapping
b) EHT: 1/2 week
- HV scan
- Correlation to EEMC
(B) Ramp-up prior to physics mode running: February 22, 2013 - March 02, 2013
- Continue with above list if not finished
- Update FGT specific tasks for shift crew vs. RHIC operation
- Exercise FGT tasks together with shift crew
Goal: Hand FGT tasks to shift crew with beginning of physics mode running
OFFLINE CREW: Anselm et al.
(4) Data taking:
Data: March 01, 2012 - June 30, 2012
On-call expert: Xuan, Bernd (Permanent on LI during Run 13)
Gas: Gerrit
DAQ: Gerard
Slowcontrol: Paul
(A) QA: Devika (Graduate student at Temple Univ.)
* Source: J-Plots / Will's plots / elog
* Record run-time info. / QA information in pre-defined sheets
* Track of live APV chips / channels vs. time
* Track of online stability, e.g pedestal stability / status
* Provide weekly status reports
(B) OFFLINE / Data processing: Anselm et al., in particular Akio
* Production of mudst's file from DAQ files: Devika
* Analysis on stability of system and various characteristic FGT quantities with tracking incl. correlation to TPC / EEMC with regular status updates
(C) Trigger:
* Trigger: MinBias and EHT
* L2: L2W
* Trigger monitoring related to W program: L0 / L2
(D) Monitoring of other spin related quantities:
* Spin monitoring: Daniel (Graduate student at Temple Univ.), TBD in SPIN PWG
(E) Reporting:
* Weekly on-call expert attends every 10:00AM meeting and provides if necessary updates / Attend 02:30PM CAD scheduling meetings
* Weekly on-call expert provides brief weekly reports in FGT meeting
* QA: Weekly reports in FGT meeting by Devika
Groups:
- surrow's blog
- Login or register to post comments